OpenStudy (anonymous):

how do i change 12 1/2% into a fraction?

OpenStudy (anonymous):

= 12.5 / 100 = 125/1000 1000 / 125 = 8 so fraction is 1/8

OpenStudy (anonymous):

to change a percentage to a fraction divide by 100 and simplify if possible